# Theory and Interpretability of Quantum Extreme Learning Machines: a Pauli-transfer matrix approach

### v1

@@ IV.4 Statistical decodability of input data
We focused above on the Pauli features $\phi_{j}$, which are generated by the encoding and transform linearly under the action of the QELM. However, they are themselves nonlinear functions of the classical input $u$. For example, with amplitude encoding (8), a single Pauli-$Z$ generates a linear readout $\langle Z\rangle\propto u$, a single Pauli-$X$ can introduce significant nonlinearities since $\langle X\rangle=\sqrt{1-(2u-1)^{2}}=1-\frac{1}{2}(2u-1)^{2}-\frac{1}{8}(2u-1)^{4}+\mathcal{O}(u^{6})$ (cf. Table 1). It is therefore important to understand the actual nonlinearities in $u$ present at the readout. This is a classical problem in reservoir computing, studied initially as memory capacity [80] and later generalized towards information processing capacity [52, 53, 81, 44]. We discuss this, as well as an alternative approach, next.
@@ IV.4.1 Nonlinear degree at readout
In order to estimate the degree of nonlinearity in $u$ available at the readout, one typically calculates the empirical correlation between the readout features $\mathbf{F}(\mathbf{u})$ [see Eq. 16] and certain nonlinear target functions $y(\mathbf{u})$. Raw monomials like $\{u^{j}\}$ are strongly correlated under typical distributions of $u$ and thus do not allow for clear identification of the nonlinear degree within this statistical approach. Following [52, 44], we therefore use orthogonal polynomials $p_{j}(u_{\alpha})$ (degree $j$) to construct orthogonal degree-$k$ targets as $y_{k}(\mathbf{u})=\prod_{i=1}^{n}p_{e_{i}}(\mathbf{u})$, with $\sum_{i=1}^{n}e_{i}=k$. Specifically, for amplitude encoding, we use (shifted) Legendre polynomials, which are orthonormal w.r.t. the uniform distribution [1,1][-1,1] or $[0,1]$; for angle encoding, we use Hermite polynomials, which are orthonormal w.r.t. the Gaussian distribution $\mathcal{N}(0,1)$.
@@
Figure 5: Ideal nonlinear capacity score $\mathcal{R}^{2}(k)$ [Eq. 43] for amplitude and rotational encodings [Eqs. 8 and 10], indicating the ability of an ideal QELM $f(\mathbf{u})=\mathbf{w}^{\top}\boldsymbol{\upphi}(\mathbf{u})$ to generate degree-$k$ monomials of the input $u_{\alpha}$ ($\alpha=1,\ldots,n$). Unless indicated, only Pauli-$Z$ features are measured in (a). Nonlinear capacity is significantly enhanced when measuring over all Paulis (b). This is also demonstrated by the score for amplitude encoding with Pauli-$X$ readout included in (a). The dash-dotted curve represents the theoretical score for $\mathcal{R}^{2}_{Z}(k)$ with amplitude encoding and $n=5$ input qubits [Eq. 44], which agrees exactly with the numerical results. $\Sigma$ denotes the integrated capacity over all $k$. The curves drop with increasing $k$ is because the quantum model can only generate square-free monomials of the Pauli features $\phi_{j}$.
@@
$\mathcal{R}^{2}(k)=1-\frac{\mathbb{E}_{\mathbf{u}}(y_{k}(\mathbf{u})-\hat{f}(\mathbf{u}))^{2}}{\mathrm{Var}_{\mathbf{u}}y_{k}(\mathbf{u})},$ (43)
to indicate whether the trained predictor $\hat{f}(\mathbf{u})$ can generate nonlinearities of degree $k$ in the input $u$  [^4] In Eq. 43, $\mathrm{Var}_{\mathbf{u}}y(\mathbf{u})$ denotes the variance of $y(\mathbf{u})$ over the distribution of $\mathbf{u}$.. $\mathcal{R}^{2}$ generalizes the geometric score (40) by taking statistical correlations into account (see Appendix C). In fact, for whitened orthonormal targets, one has $\mathcal{R}^{2}(k)=\gamma_{k}^{2}$. In the case of amplitude encoding (8) with Pauli-$Z$ readout, the number of representable degree-$k$ targets is $\binom{n}{k}$, which exponentially decreases compared to the total number of degree-$k$ monomials $\binom{n+k-1}{k}$, giving an expected score
$\mathcal{R}^{2}_{Z}(k)=\begin{cases}\binom{n}{k}/\binom{n+k-1}{k}&\text{for }k\leq n\\ 0&\text{for }k>n\end{cases}$ (44)
In Fig. 5, we first display the nonlinear capacity that can be ideally achieved from the encoded state without effects of feature mixing, i.e., we take $\mathbf{F}=\boldsymbol{\upphi}$ in (25). As shown in panel (a), a Pauli-$X$ readout (black symbols with dotted curves) generates even degree monomials, behaving similarly to angle encoding with Pauli-$Z$ observables (connected open symbols). In Fig. 5(b) we show the nonlinear capacity score $\mathcal{R}^{2}(k)$ obtained for a complete Pauli observable set. Notably, the presence of non-$Z$ Paulis significantly enhances the nonlinearity present in the readout, as becomes clear by comparison to the dash-dotted curve corresponding to $\mathcal{R}^{2}_{Z}(k)$ for $n=5$ [Eq. 44]. Moreover, angle and amplitude encodings show a similar behavior. The decay of $\mathcal{R}^{2}(k)$ with $k$ stems from the fact that the encodings used here can only generate square-free monomials of the Pauli features $\phi_{j}$, while we test against (the practically relevant case of) all possible degree-$k$ targets.
@@ IV.4.2 Nonlinear monomials at readout
Figure 8: Ability to construct monomials of the input $\{x,y,z\}$ for amplitude encoding (8) and rotational encoding (10) with $n=D=3$. (a) Decodability score $\kappa_{j}^{2}$ according to Eq. 45. (b) Projector matrix $A^{+}A$, where nonzero off-diagonals correspond to $\kappa_{j}^{2}<1$ and indicate decoding is possible only within a certain accuracy. The underlying Taylor expansion is performed up to order $r=4$ in each $u_{j}=\{x,y,z\}$ and changes are mild upon further increasing $r$. The scores and matrices are averaged over a uniform grid of expansion points $\mathbf{u}_{0}\in[0.1,0.9]^{3}$ for amplitude encoding and $\mathbf{u}_{0}\in[-1,1]^{3}$ for rotational encoding. For readability, monomial labels are incompletely displayed in (a) and replaced by column/row indices in (b).
Raw monomials in $u_{\alpha}$ are often relevant features for practical applications, such as autoregressive learning of dynamical systems. Let thus $\mathcal{B}(\mathbf{u})=\{1,u_{1},\ldots,u_{1}u_{2},\ldots,\mathcal{O}(u^{r})\}$ be the vector of size $b=\binom{D+r}{r}$ of monomials of $\{u_{\alpha}\}_{1\leq\alpha\leq D}$ up to degree $r$. The order-$r$ Taylor expansion of the Pauli features $\{\phi_{j}(\mathbf{u})\}_{1\leq j\leq q}$ within a small region around some operating point $\mathbf{u}_{0}$ (e.g., data mean) can then be written as $\boldsymbol{\upphi}(\mathbf{u})=A\mathcal{B}(\mathbf{u})$ with a coefficient matrix $A\in\mathbb{R}^{q\times b}$. As before, we can define a geometric decodability score
@@
which measures how accurately a certain monomial $\mathpzc{b}\in\mathcal{B}$ can be constructed from linear combinations of $\phi_{j}$. This score depends on the operating point $\mathbf{u}_{0}$, but is independent of quantum properties of the reservoir, and ranges between 0 and 1 (no/exact construction possible). Intermediate values indicate that construction is possible only within an error $\sim\sqrt{1-\kappa_{j}^{2}}$ due to off-diagonal terms in the $j$-th row/column of the projector $A^{+}A$. A somewhat artificial dependence of $\kappa_{j}^{2}$ on the Taylor order $r$ (owing to the rank constraint) can be mitigated by considering large $r$ and $b$. Indeed, most of the $\kappa_{j}^{2}$ decrease upon increasing $r$, except for a few stable ones. These include monomials that arise directly from the tensor product structure of the encoding state. For amplitude encoding (8), these are all square-free monomials up to order $n$ (e.g., for $n=D=3$: $\mathcal{B}_{\text{stable}}=\{1,x,y,z,xy,xz,yz,xyz\}$).
The typical behavior of $\kappa_{j}^{2}$ and of the projectors $A^{+}A$ (averaged over a range of expansion points) is illustrated in Fig. 8. While the score typically decays with increasing nonlinear degree, rotational encoding tends to have a greater expressivity towards higher-order monomials. This is associated with less leakage into off-diagonal terms of $A^{+}A$ [see Fig. 8(b)], and can explain why rotational encoding often leads to robust performance of QELMs.
IV.5 Summary
Feature decodability is the underlying concept that determines the expressivity of a QELM. Pauli features are the fundamental “building blocks” of the QELM expressivity. Any other nonlinear feature required for the learning task must be constructible from Pauli features. For the encoding types considered here, expressing simple monomials is often only possible up to a certain approximation error ($\kappa_{j}^{2}<1$), which increases with the size of the data domain. Knowing the structure of the readout features allows to design feature engineering strategies to enhance the expressivity at low cost.
@@ V.1 Time series modeling with QELMs
If the dynamics is Markovian, it can be modeled solely with a NG-RC or QELM based on an instantaneous feature map [84, 29]. Accordingly, we formulate the training problem of the QELM (25) as
$\text{QELM:}\qquad\mathbf{u}_{t+1}=\mathbf{f}(\mathbf{u}_{t})=W^{\top}R\boldsymbol{\upphi}(\mathbf{u})$ (48)
@@
$\boldsymbol{\upphi}_{\text{amp}}^{n=3}(\mathbf{u})=\Bigl\{$ $1,2\sqrt{(1-x)x},2x-1,2\sqrt{(1-y)y},2(2x-1)\sqrt{-((y-1)y)},4\sqrt{(x-1)x(y-1)y},2y-1,$ (50)
$2\sqrt{-((x-1)x)}(2y-1),(2x-1)(2y-1),2\sqrt{(1-z)z},2(2x-1)\sqrt{-((z-1)z)},$ (51)
$2(2y-1)\sqrt{-((z-1)z)},2(2x-1)(2y-1)\sqrt{-((z-1)z)},4\sqrt{(x-1)x(z-1)z},$ (52)
@@
$8\sqrt{-((x-1)x(y-1)y(z-1)z)},2z-1,2\sqrt{-((x-1)x)}(2z-1),(2x-1)(2z-1),$ (54)
$2\sqrt{-((y-1)y)}(2z-1),2(2x-1)\sqrt{-((y-1)y)}(2z-1),4(2z-1)\sqrt{(x-1)x(y-1)y},$ (55)
$(2y-1)(2z-1),2\sqrt{-((x-1)x)}(2y-1)(2z-1),(2x-1)(2y-1)(2z-1)\Bigr\}$ (56)
We will also consider the amplitude encoding scheme of Eq. 9, which is obtained by replacing in Eq. 56 $\{x,y,z\}\to\{x^{2},y^{2},z^{2}\}$. Depending on the dynamical system, we will observe differences in the performance.
In order to confine the trajectory to the interval $\mathbf{u}\in[0,1]^{3}$ or $1,1]^{3}$, as required by the encoding, we apply the transformation $\mathbf{u}\to\mathbf{u}^{\prime}=\boldsymbol{\alpha}\cdot(\mathbf{u}-\mathbf{m})$ with $\boldsymbol{\alpha},\mathbf{m}\in\mathbb{R}^{3}$ to the state. This leads to the rescaled Lorenz-63 model:
@@
This vector field-level transformation allows us to directly compare the learned coefficients with the true ones.

### v2

@@ IV.4 Statistical decodability of input data
We focused above on the Pauli features $\phi_{j}$, which are generated by the encoding and transform linearly under the action of the QELM. However, they are themselves nonlinear functions of the classical input $\mathbf{u}\in\mathbb{R}^{D}$. For example, with amplitude encoding (8), a single Pauli-$Z$ generates a linear readout $\langle Z\rangle\propto u$, a single Pauli-$X$ can introduce significant nonlinearities since $\langle X\rangle=\sqrt{1-(2u-1)^{2}}=1-\frac{1}{2}(2u-1)^{2}-\frac{1}{8}(2u-1)^{4}+\ldots$ (cf. Table 1). It is therefore important to understand the actual nonlinearities in $u$ present at the readout. This is a classical problem in reservoir computing, studied initially as memory capacity [90] and later generalized in terms of information processing capacity [54, 91, 55, 47]. We discuss this, as well as an alternative approach, next.
@@ IV.4.1 Nonlinear degree at readout
In order to estimate the degree of nonlinearity in $u$ available at the readout, one typically calculates the empirical correlation between the readout features $\mathbf{F}(\mathbf{u})$ [see Eq. 16] and certain nonlinear target functions $y(\mathbf{u})$. Raw monomials like $\{u^{j}\}$ are strongly correlated under typical distributions of $u$ and thus do not allow for clear identification of the nonlinear degree within this statistical approach. Following [54, 47], we therefore use orthogonal polynomials $p_{j}(u_{\alpha})$ (degree $j$) to construct orthogonal degree-$k$ targets as $y_{\mathbf{e}}(\mathbf{u})=\prod_{\alpha=1}^{D}p_{e_{\alpha}}(u_{\alpha})$, with $\sum_{\alpha=1}^{D}e_{\alpha}=k$. Specifically, for amplitude encoding, we use (shifted) Legendre polynomials, which are orthonormal w.r.t. the uniform distribution [1,1][-1,1] or $[0,1]$; for angle encoding, we use Hermite polynomials, which are orthonormal w.r.t. the Gaussian distribution $\mathcal{N}(0,1)$  [^5] Due to periodicity of rotational encoding [Eq. 10], the achievable capacity can be slightly reduced when sampling $u\sim\mathcal{N}$..
@@
Figure 5: Ideal nonlinear capacity score $\mathcal{R}^{2}(k)$ [Eq. 44] for amplitude and rotational encodings [Eqs. 8 and 10], indicating the ability of an ideal QELM $f(\mathbf{u})=\mathbf{w}^{\top}\boldsymbol{\upphi}(\mathbf{u})$ to generate degree-$k$ monomials of the input $u_{\alpha}$ ($\alpha=1,\ldots,D$ and $n=D$). Unless indicated, only Pauli-$Z$ features are measured in (a). Nonlinear capacity is significantly enhanced when measuring over all Paulis (b). This is also demonstrated by the score for amplitude encoding with Pauli-$X$ readout included in (a). The dash-dotted curve represents the theoretical score for $\mathcal{R}^{2}_{Z}(k)$ with amplitude encoding and $n=5$ input qubits [Eq. 45], which agrees exactly with the numerical results. $\Sigma$ denotes the integrated capacity over all $k$.
@@
$\mathcal{R}^{2}(k)=\frac{1}{N_{k}}\sum_{|\mathbf{e}|=k}\mathcal{R}^{2}[y_{\mathbf{e}}],\qquad\mathcal{R}^{2}[y_{\mathbf{e}}]=1-\frac{\mathbb{E}_{\mathbf{u}}(y_{\mathbf{e}}(\mathbf{u})-\hat{f}(\mathbf{u}))^{2}}{\mathrm{Var}_{\mathbf{u}}y_{\mathbf{e}}(\mathbf{u})},$ (44)
averaged over all $N_{k}=\binom{D+k-1}{k}$ degree-$k$ monomials, to indicate whether the trained predictor $\hat{f}(\mathbf{u})$ can generate nonlinearities of degree $k$ in the $(D=n)$-dimensional input  [^6] In Eq. 44, $\mathrm{Var}_{\mathbf{u}}y(\mathbf{u})$ denotes the variance of $y(\mathbf{u})$ over the distribution of $\mathbf{u}$.. $\mathcal{R}^{2}$ generalizes the geometric score (41) by taking statistical correlations into account (see Appendix C). In the special case where the targets are whitened orthonormal features, one has $\mathcal{R}^{2}[y_{r}]=\gamma_{r}^{2}$. In the case of amplitude encoding (8) with Pauli-$Z$ readout, the number of representable degree-$k$ targets is $\binom{D}{k}$, giving an averaged score
$\mathcal{R}^{2}_{Z}(k)=\begin{cases}\binom{D}{k}/\binom{D+k-1}{k}&\text{for }k\leq D,\\ 0&\text{for }k>D\end{cases}$ (45)
which exponentially decreases with $k$.
In Fig. 5, we first display the nonlinear capacity that can be ideally achieved from the encoded state without effects of feature mixing, i.e., we take $\mathbf{F}=\boldsymbol{\upphi}$ in (25). As shown in panel (a), a Pauli-$X$ readout (black symbols with dotted curves) generates even degree monomials, behaving similarly to angle encoding with Pauli-$Z$ observables (connected open symbols). In Fig. 5(b) we show the nonlinear capacity score $\mathcal{R}^{2}(k)$ obtained for a complete Pauli observable set. Notably, the presence of non-$Z$ Paulis significantly enhances the nonlinearity present in the readout, as becomes clear by comparison to the dash-dotted curve corresponding to $\mathcal{R}^{2}_{Z}(k)$ for $D=n=5$ [Eq. 45]. Moreover, angle and amplitude encodings show a similar behavior. We attribute the decay of $\mathcal{R}^{2}(k)$ with $k$ mainly to fact that the product state encoding only generates square-free monomials of the Pauli features $\phi_{j}$, while we test against (the practically relevant case of) all possible degree-$k$ targets  [^7] This argument is not fully rigorous, since features like $\phi_{x}^{(i)}(u_{i})$ have infinite Taylor expansion in $u_{i}$ and can thus correlate locally with raw monomials $u_{i}^{m}$..
@@ IV.4.2 Nonlinear monomials at readout
Figure 8: Ability to construct monomials of the input $\{x,y,z\}$ for amplitude encoding (8) and rotational encoding (10) with $n=D=3$. (a) Decodability score $\kappa_{j}^{2}$ according to Eq. 46. (b) Projector matrix $A^{+}A$, where nonzero off-diagonals correspond to $\kappa_{j}^{2}<1$ and indicate decoding is possible only within a certain accuracy. The underlying Taylor expansion is performed up to order $r=4$ in each $u_{j}=\{x,y,z\}$ and changes are mild upon further increasing $r$. The scores and matrices are averaged over a uniform grid of expansion points $\mathbf{u}_{0}\in[0.1,0.9]^{3}$ for amplitude encoding (with some margin to the singular boundaries) and $\mathbf{u}_{0}\in[0,1]^{3}$ for rotational encoding. For readability, monomial labels are incompletely displayed in (a) and replaced by column/row indices in (b).
Raw monomials in $u_{\alpha}$ are often relevant features for practical applications, such as autoregressive learning of dynamical systems. Let thus $\mathcal{B}(\mathbf{u})=\{1,u_{1},\ldots,u_{1}u_{2},\ldots,\prod_{\alpha}u_{\alpha}^{r}\}$ be the vector of size $b=(r+1)^{D}$ of monomials of $\{u_{\alpha}\}_{1\leq\alpha\leq D}$ up to degree $r$ (per coordinate). The order-$r$ Taylor expansion of the Pauli features $\{\phi_{j}(\mathbf{u})\}_{1\leq j\leq q}$ within a small region around some operating point $\mathbf{u}_{0}$ (e.g., data mean) can then be written as $\boldsymbol{\upphi}(\mathbf{u})\simeq A(\mathbf{u}_{0})\mathcal{B}(\mathbf{u}-\mathbf{u}_{0})$ with a coefficient matrix $A(\mathbf{u}_{0})\in\mathbb{R}^{q\times b}$. As before, we can define a geometric decodability score
@@
which measures how accurately a certain monomial $\mathpzc{b}\in\mathcal{B}$ can be constructed from linear combinations of $\phi_{j}$. This score depends on the operating point $\mathbf{u}_{0}$, but is independent of quantum properties of the reservoir, and ranges between 0 and 1 (no/exact construction possible). Intermediate values indicate that construction is possible only within an error $\sim(1-\kappa_{j}^{2})^{1/2}$ due to off-diagonal terms in the $j$-th row/column of the projector $A^{+}A$. A somewhat artificial dependence of $\kappa_{j}^{2}$ on the Taylor order $r$ (owing to the rank constraint) can be mitigated by considering large $r$ and $b$. Indeed, most of the $\kappa_{j}^{2}$ decrease upon increasing $r$, except for a few stable ones. These include monomials that arise directly from the tensor product structure of the encoding state. For amplitude encoding (8) including all $Z$-strings, these are all square-free monomials up to order $n$ (e.g., for $n=D=3$: $\mathcal{B}_{\text{stable}}=\{1,x,y,z,xy,xz,yz,xyz\}$).
The typical behavior of $\kappa_{j}^{2}$ and of the projectors $A^{+}A$ (averaged over a range of expansion points) is illustrated in Fig. 8. While the score typically decays with increasing nonlinear degree, rotational encoding tends to have a greater expressivity towards higher-order monomials. This is associated with less leakage into off-diagonal terms of $A^{+}A$ [see Fig. 8(b)], and contributes to the robust performance often observed for rotational encoding.
IV.5 Summary and Discussion
It is useful to briefly discuss the injectivity property of the feature map [49], which ensures that distinct inputs map to distinct readouts. A standard condition for local injectivity of $\mathbf{F}(\mathbf{u})=R\boldsymbol{\upphi}(\mathbf{u})$ [Eq. 25] is a full-rank Jacobian:
$\mathrm{rank}\,RJ_{\boldsymbol{\upphi}}(\mathbf{u})=D,\qquad J_{\boldsymbol{\upphi}}(\mathbf{u})=\frac{\partial\boldsymbol{\upphi}(\mathbf{u})}{\partial\mathbf{u}}\in\mathbb{R}^{{q}\times D}.$ (47)
For the considered product state encodings it is straightforward to check that $\boldsymbol{\upphi}(\mathbf{u})$ is injective on the respective domains in $\mathbb{R}^{D}$ (possibly after removing one endpoint). Hence, local injectivity of $\mathbf{F}(\mathbf{u})$ is determined by whether $R$ is injective on the tangent space $\operatorname{range}\,J_{\boldsymbol{\upphi}}(\mathbf{u})$, i.e. whether $\ker R\cap\operatorname{im}J_{\boldsymbol{\upphi}}(\mathbf{u})=\{0\}$. In a generic setting, this condition is satisfied (at fixed $\mathbf{u}$) if $\mathrm{rank}\,R\geq D$, which is the case for adequate measurement schemes and sufficiently scrambling dynamics.
Decodability of Pauli features can be understood as the underlying concept that determines the expressivity of a QELM with initial-state encoding and data-independent reservoir dynamics. High decodability scores are associated with high-rank effective PTMs $R$, which are achieved for random unitaries. Strongly mixing quantum channels can help expressivity and nonlinear processing in QELMs [44, 14, 50], and are thus often beneficial for training performance apart from concentration effects [53]. By contrast, QRCs with internal memory states often work best at a point where random matrix behavior just begins to dominate (“edge of chaos”) (see, e.g., [41, 95, 96]). An alternative to increasing randomness is to devise task-aligned reservoir structures and measurement schemes [2, 43, 97, 50, 32]. Knowing the connection between the readout and the encoding-generated features via the PTM framework can thus inform feature engineering strategies to enhance expressivity.
@@ V.1 Time series modeling with QELMs
If the dynamics is Markovian, it can be modeled solely with a NG-RC or QELM based on an instantaneous feature map (see [32] for examples). Accordingly, we formulate the training problem of the QELM (25) as
$\text{QELM:}\qquad\mathbf{u}_{t+1}=\mathbf{f}(\mathbf{u}_{t})=W^{\top}R\boldsymbol{\upphi}(\mathbf{u}_{t})$ (50)
@@
$\boldsymbol{\upphi}_{\text{amp}}^{n=3}(\mathbf{u})=\Bigl\{$ $1,2\sqrt{(1-x)x},2x-1,2\sqrt{(1-y)y},2(2x-1)\sqrt{-((y-1)y)},4\sqrt{(x-1)x(y-1)y},2y-1,$
$2\sqrt{-((x-1)x)}(2y-1),(2x-1)(2y-1),2\sqrt{(1-z)z},2(2x-1)\sqrt{-((z-1)z)},$
$2(2y-1)\sqrt{-((z-1)z)},2(2x-1)(2y-1)\sqrt{-((z-1)z)},4\sqrt{(x-1)x(z-1)z},$
@@
$8\sqrt{-((x-1)x(y-1)y(z-1)z)},2z-1,2\sqrt{-((x-1)x)}(2z-1),(2x-1)(2z-1),$
$2\sqrt{-((y-1)y)}(2z-1),2(2x-1)\sqrt{-((y-1)y)}(2z-1),4(2z-1)\sqrt{(x-1)x(y-1)y},$
$(2y-1)(2z-1),2\sqrt{-((x-1)x)}(2y-1)(2z-1),(2x-1)(2y-1)(2z-1)\Bigr\}$
We will also consider the amplitude encoding scheme (9), which has a similar polynomial structure.
In order to confine the trajectory to the interval $\mathbf{u}\in[0,1]^{3}$ or $1,1]^{3}$, as required by the encoding, we apply the transformation $\mathbf{u}\to\mathbf{u}^{\prime}=\boldsymbol{\alpha}\odot(\mathbf{u}-\mathbf{m})$ with $\boldsymbol{\alpha},\mathbf{m}\in\mathbb{R}^{3}$ to the state (we typically use the same $\alpha$ for all coordinates). This leads to the rescaled Lorenz-63 model:
@@
This vector field-level transformation simplifies comparison between the learned coefficients and the true ones.